THE FAR INFRARED EXPERIMENTAL FACILITY AT THE CANADIAN LIGHT SOURCE

\underline{D.R.T APPADOO} and T. MAY,, Canadian Light Source Inc., Saskatoon, Saskatchewan S7N 0X4, Canada,; and A.R.W. MCKELLAR, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario K1A 0R6, Canada.

The far-infrared (far-IR) experimental facility at the Canadian Light Source Inc. (CLSI) has been dedicated primarily to high-resolution gas-phase spectroscopic studies of stable and unstable molecule. Using long wavelength optics, the Bending Magnet’s infrared radiation is steered to a Bruker IFS125 HR spectrometer for ultra-high spectral resolution (≥ 0.0009 cm^{−1}, MOPD = 9.4 m) gas-phase studies. This report describes our efforts at the CLSI in getting the far-IR experimental facility ready for users. In particular, we will compare the brightness and noise level of the IR synchrotron radiation at the CLSI to that of a globar in the far-IR region; we will also study the lineshapes and linewidths.